OpenStudy (anonymous):

I am having some problems solving a couple different equations. Please help. The first one says solve by completing the square: x^2+3x-70=0

OpenStudy (cwrw238):

the aim of completing the square is to obtain a perfect square on one side of equation and a number on other side

OpenStudy (anonymous):

since you said to complete the square move the -70 over to the right side of the = sign x^2+3x=70 now take half of the coefficient of x and square it... (3/2)^2. this is what you need to make the left side a perfect square trinomial. and don't forget to add the same amount to the right. x^2+3x+(3/2)^2=70+(3/2)^2 now the left side is factored as (x+3/2)^2=70+9/4 can you take it from here?

OpenStudy (anonymous):

I have tried that on another problem that i had and the answers I came up with were wrong. So I have an error in one of my steps. Just not sure where it is.

OpenStudy (anonymous):

I can sure try to take it from there.

OpenStudy (anonymous):

btw... your answer should be -10, 7. solving the equation by factoring should be much easier...

OpenStudy (anonymous):

Ok I am not getting close to that. I don't know where I am going wrong.

OpenStudy (anonymous):

I'm confused that you mention about different equation, anyway here is how to complete square: x^2+3x-70=0 (x + 3/2)² - 70 - 9/4 (x + 3/2)² - 289/4

OpenStudy (anonymous):

Different problems I am working on...this is just one of them I needed help with.

OpenStudy (anonymous):

Ok thats what I did....so you subtract the 289/4?

OpenStudy (anonymous):

That's called complete square! finished!

OpenStudy (anonymous):

so how do I come up with the answer? -10 and 7, which was correct in mymathlab.

OpenStudy (anonymous):

What you ask now is solve it by fraction!

OpenStudy (anonymous):

Sorry fractions are something that I don't understand how to do especially when using in formulas.

OpenStudy (anonymous):

When you apply product and sum method to factor: (x + 10) ( x -7) = 0

OpenStudy (anonymous):

x + 10 = 0 --> x = -10 x - 7 = 0 --> x = 7

OpenStudy (anonymous):

ahha...got it...thanks for explaining further.

OpenStudy (anonymous):

:)